OpenStudy (anonymous):

What is the 6th root of 256? ...left in radical notation

OpenStudy (anonymous):

6th root of 256 = 2.5198421

OpenStudy (anonymous):

thanks for responding, but I was looking for the answer in radical notation

OpenStudy (accessdenied):

What I would do is break it into prime factorization and try to take out "perfect 6th root" factors.

OpenStudy (pfenn1):

\[\sqrt[6]{256}\]

OpenStudy (anonymous):

ok, \[\sqrt[6]{256} = \sqrt[6]{64\times4} = 2\sqrt[6]{4} ?\]

OpenStudy (accessdenied):

Yeah, that's what I get as well. :)

OpenStudy (anonymous):

Hmm...the answer Im suppposed to get is \[2\sqrt[3]{2}\]

OpenStudy (accessdenied):

Ohh, I see that we should also use the fact that "(2^2)^(1/6) = 2^(2/6) = 2^(1/3)"

OpenStudy (accessdenied):

or in radical form... \( \sqrt[6]{2^2} = (\sqrt[6]{2})^2 = \sqrt[3]{2} \)

OpenStudy (anonymous):

wow, thanks for your response...I guess when they mean "write the expression in simplest form" they really mean simplest form!

OpenStudy (anonymous):

I don't always "see" the simplest form in these types of problems!

OpenStudy (accessdenied):

You're welcome. :) I was thinking about if we needed the radical reduced as well in this case since I've never ran across that situation in my class. :P

OpenStudy (anonymous):

Do you think since \[2\sqrt[6]{4} = 2\sqrt[3]{2}\] it is safe to assume if there was anothther perfect sqaure inside the square root sign, where the 4 is, then I would have to simplify further?

OpenStudy (accessdenied):

Yeah. I think it's easiest to understand the simplification when you express the radical as a rational power, a fraction. If we have some expression to a power inside the radical, then we can consider reducing the power and index as if they were a fraction. \(\sqrt[6]{25} = \sqrt[6]{5^2} = 5^{2/6} = 5^{1/3} \) A situation I'm unsure of, however, is if you have an extra factor there, though, like 3*25. It doesn't really seem simpler to break it into TWO radicals to me. That may be something to ask the teacher about. :P

OpenStudy (anonymous):

thanks for your insight! It has helped me see better when to simplify further

OpenStudy (accessdenied):

You're welcome! Again, these are some really rare cases to deal with, from my experience. :P

OpenStudy (anonymous):

I need to post another question, I have to close this one!